OpenStudy (anonymous):

Which system of linear equations has no solution? A. 2x − y = −1 2x + y = 2 B. 4x − y = 3 3x − 2y = 3 C. −2x + 4y = −16 4x − 2y = −1 D. 2x − y = −1 −2x + y = −1

OpenStudy (anonymous):

Just go through each one and solve the systems until one doesnt work, so, for A, 2x - y = -1 + 2x + y = 2

OpenStudy (anonymous):

they y's cancel out because you have -1y + 1y which equals 0 so youre left with 4x = 1

OpenStudy (anonymous):

solve that, and x = 1/4

OpenStudy (anonymous):

plug that value of x into either one of the original equations and then solve for y, so. 2(1/4) - y = -1. 0.5 - y = -1 -y = -1.5 y=1.5

OpenStudy (anonymous):

that one has a solution of (1/4, 3/2). do the same thing for the others until one doesnt work

OpenStudy (anonymous):

*hint - if your values dont conveniently cancel out the way they do with A, you might have to multiply them until they do cancel out by numbers like -1, or 2 or so on*

OpenStudy (anonymous):

So your saying that A is the answer

OpenStudy (anonymous):

No, im saying that A is not an answer because it does have a solution

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

Which system of linear equations shares the same line? A. 6x − 3y = 15 8x − 4y = 20 B. 6x − 2y = 2 x + y = C. 3x − 3y = 15 9x − 6y = 6 D. −4x + 4y = −4 2x + 2y = 2

OpenStudy (anonymous):

what is the end of B?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

Oh this is the last question

OpenStudy (anonymous):

Try to convert each into y=mx+b form to make it easier. A. -3y = -6x + 15 divide both sides by -3 y = 2x - 5 8x - 4y = 20. Subtract the 8x from both sides. -4y = -8x + 20. Divide by -4 on both sides. y = 2x - 5. the two are the same so they would make the same line, A is your answer

OpenStudy (anonymous):

I got two wrong